### v3

# SemioLLM: Evaluating Large Language Models for Diagnostic Reasoning from Unstructured Clinical Narratives in Epilepsy

Abstract
Large Language Models (LLMs) have been shown to encode clinical knowledge. Many evaluations, however, rely on structured question-answer benchmarks, overlooking critical challenges of interpreting and reasoning about unstructured clinical narratives in real-world settings. In this study we task eight Large Language models including two medical models (GPT-3.5, GPT-4, Mixtral-8x7B, Qwen-72B, LlaMa2, LlaMa3, OpenBioLLM, Med42) with a core diagnostic task in epilepsy: mapping seizure description phrases—after targeted filtering and standardization—to one of seven possible seizure onset zones using likelihood estimates. Most models yield results that often match the ground truth and even approach clinician-level performance after prompt engineering. Specifically, clinician-guided chain-of-thought reasoning leading to the most consistent improvements. Performance was further strongly modulated by clinical in-context impersonation, narrative length and language context (13.7%, 32.7% and 14.2% performance variation, respectively). However, expert analysis of reasoning outputs revealed that correct prediction can be based on hallucinated knowledge and inaccurate source citation, underscoring the need to improve interpretability of LLMs in clinical use. Overall, SemioLLM provides a scalable, domain-adaptable framework for evaluating LLMs in clinical disciplines where unstructured verbal descriptions encode diagnostic information. By identifying both the strengths and limitations of LLMs, our work contributes to testing the applicability of foundational AI systems for healthcare.
1 Introduction
Large Language Models (LLMs) have shown significant potential in leveraging clinical knowledge on structured question answering (Q&A) datasets such as MedQA (Jin et al., 2021), PubMedQA (Jin et al., 2019), MedMCQA (Pal et al., 2022) and BioASQ-QA (Krithara et al., 2023) in multiple medical domains (Sarvari et al., 2024; Savage et al., 2024; Singhal et al., 2023a, b; van Diessen et al., 2024). While Q&A set-ups are advantageous as they provide a clear ground-truth for model testing, they oversimplify clinical decision-making (Hager et al., 2024), which often relies on extracting crucial diagnostic information from unstructured patient interviews containing complex, irrelevant and everyday language (Alaa et al., 2025; ESR, 2023; Raji et al., 2025). While LLMs have shown remarkable capabilities in extracting meaningful information from unstructured text for a wide range of downstream tasks in other domains (Bommasani et al., 2021), their ability to do so in clinical contexts remains poorly understood. Thus, clinical narratives from patients provide crucial diagnostic information to be leveraged by LLMs especially when structured input is limited or absent (van Diessen et al., 2024; Yan et al., 2024). However, it is an open and largely unexplored question of how well LLMs can extract and interpret clinically meaningful information from unstructured clinical narratives to support real-world diagnostic reasoning.
Neurological disorders, such as epilepsy, provide a particularly compelling use case to explore this question, as behavioral and sensory symptoms can often be directly linked to underlying brain pathologies. In epilepsy, clinicians routinely base clinical decisions on patient and witness accounts of seizure manifestations—known as semiology (Nowacki & Jirsch, 2017; Tufenkjian & Lüders, 2012). Especially during early diagnostic evaluations, correctly interpreting seizure symptoms is crucial for guiding follow-up procedures such as brain imaging, EEG and surgical planning (Beniczky et al., 2022; Lüders et al., 2006). Seizure descriptions contain diagnostically relevant information about the seizure origin in the brain, and help classifying seizure types and syndromes. For example, repetitive chewing, swallowing, or lip-smacking strongly indicate temporal lobe involvement (Wiebe et al., 2001), while excessive limb movements or pelvic thrusting are indicative of frontal lobe seizures (Lüders et al., 2006). Accurate localization of the seizure onset zone (SOZ) is particularly important for patients with drug-resistant epilepsy, where surgical resection of the SOZ remains the only potentially curative treatment option (Sisodiya & Goldstein, 2007; Wiebe et al., 2001).
In this study, we develop a structured and automated evaluation framework, SemioLLM, that benchmarks LLMs’ ability to extract and translate diagnostically relevant information from seizure descriptions into probabilistic seizure locations in the brain. Using an annotated database linking over 1,200 seizure descriptions to seizure foci (Alim-Marvasti et al., 2022), we evaluate eight LLMs, including proprietary and open-source models (GPT-3.5 (Brown et al., 2020), GPT-4 (Achiam et al., 2023), Mixtral-8x7B (Jiang et al., 2024), Qwen-72B (Bai et al., 2023), LlaMa2 70B (Touvron et al., 2023), and LlaMa3 70B (Dubey et al., 2024), OpenBioLLM, Med42 (Christophe et al., 2024)). We systematically analyze model accuracy, confidence, calibration and reasoning, benchmarking their outputs in comparison to evaluation by a clinical domain expert.
Our study reveals several key insights: (i) We demonstrate that LLMs can predict seizure onset zones (SOZ) based on unstructured seizure descriptions, significantly outperforming chance-level predictions. Importantly, with refined prompting techniques, for example chain-of-thought (CoT) reasoning, their accuracy substantially improves and approaches clinician-level performance. (ii) Many models exhibit reasonable trustworthiness, as assessed through an entropy-based confidence measure, where confidence similarly improves with prompt-engineering. Notably, GPT-4 and Mixtral-8x7B demonstrated an optimal balance of accuracy and confidence. (iii) Through an extensive manual reasoning assessment conducted by a domain specific clinical expert, we identify that GPT-4 demonstrates superior capabilities in integrating domain knowledge, clinical inference, and evidence verification. Mixtral-8x7B, while competitive in text comprehension, exhibits notable limitations in reasoning and accurate source citation, underscoring areas for future model refinement. (iv) Our analysis identifies key factors that influence LLM diagnostic performance in processing seizure descriptions. First, we observe a U-shaped relationship between description length and accuracy, with both very short and highly detailed narratives yielding better results than those of moderate length. Second, prompting models to impersonate clinical experts markedly improves both accuracy and confidence, suggesting better alignment with domain-specific reasoning. Third, our multilingual evaluation shows that while English-trained models perform well even when processing non-English clinical narratives, accuracy declines for different prompt languages, underscoring the need for targeted multilingual user adaptation.
In summary, our study provides a systematic and in-depth investigation of LLMs in epilepsy diagnostics based solely on verbal symptom descriptions, identifying prompt strategies and expert impersonation as the most significant factors that improved diagnostic accuracy on average by 10% and 13.8%, respectively. Unlike existing structured Q&A evaluations, SemioLLM demonstrates how LLMs can translate unstructured clinical narratives into probabilistic diagnostic decisions, and identifies conditions under which both accuracy and confidence can be improved. Our framework provides a practical guideline for similar deployments and is easily transferable to other clinical specialties where symptom descriptions inform diagnostic decisions, potentially improving early diagnosis and treatment planning for patients with complex neurological and other medical conditions.
2 Results
Our experimental pipeline of semioLLM is illustrated in Figure 1. Evaluated LLMs are given a text describing seizure symptoms taken from a public database (Alim-Marvasti et al., 2022). In the database, each seizure description is linked to one of 7 brain regions containing the respective seizure onset zone (SOZ). This ground truth was determined based on the fact that post-surgical resection of the SOZ patients remained seizure free for at least one year.
2.1 Prompt strategies significantly boost performance
We assess classification performance for SOZ localization using the F1 score, comparing LLMs against a clinical evaluation and a naive classifier (lower bound, see Methods section 1.7). We further compare the zero shot setting (ZS) to 4 different prompt strategies: In the few-shot (FS) (Brown et al., 2020) condition, we leverage in-context learning (ICL) and provide each model with representative input-output pairs of seizure descriptions and their SOZs. We employ Chain-of-thought (CoT) prompting (Wei et al., 2022), which provides step-by-step reasoning to improve the models’ ability to handle complex reasoning tasks. We also implement a task-specific strategy, FS-CoT, which combines the desired input-output mapping with a reasoning patterns curated by an expert clinician, mimicking the diagnostic reasoning of an epileptologist. Finally, we employ self-consistency (SC) (Wang et al., 2022), generating multiple reasoning paths and determining predictions through majority voting, which have been shown to enhance inference robustness and produce more reliable predictions for complex tasks.
As shown in Figure 2(a), most models perform just above chance level in the zero-shot (ZS) condition. Exceptions are Mixtral-8x7B and GPT-4, which achieved substantially higher F1 scores than all other models of 51.66% and 52.27%, respectively (95% CI: [51.43, 51.90], [52.04, 52.50]), comparable to the clinician’s performance: Clinician 1 - 48.77% (95% CI: [48.53, 49.02]) and Clinician 2 - 46.75% (95% CI: [46.51, 46.99]). Importantly, we observed a substantial performance increase when introducing prompt-engineering across all models: Median F1 improvement relative to ZS was 6.49% for Few-Shot prompting, 9.62% for CoT, 9.49% in FS-CoT, and 10.02% increase for SC. Note that GPT-4, maintained a consistently high performance in all conditions, only showing a modest gain from 52.27% in ZS to 53.44% with SC. In addition to general-purpose LLMs, we evaluated two medical-specific models—OpenBioLLM-70B and Med42-70B. OpenBioLLM-70B performed well, particularly in CoT (52.19) and SC (53.06), approaching score in SC (53.44), but showed weaker ZS performance (38.56). Med42-70B exhibited better calibration than OpenBioLLM-70B (refer Figure 2(c)), yet lower overall scores (e.g., 46.70 in CoT, 47.63 in SC). While, both medical LLMs approached some general models in specific settings, neither consistently matched the performance of the top general-purpose models such as GPT-4.0 (mean 52.10) or Mixtral-8x7B (mean 50.58) - which achieved high accuracy for seizure-onset zone prediction, even without any external guidance (ZS). However, other models are able to match this through prompt-engineering. Interestingly, FS-CoT and SC both demonstrate the highest positive impact even though relying on different strategies-with FS-CoT providing expert-curated reasoning patterns that guide model outputs, and SC by enhancing robustness through the aggregation of multiple independent reasoning paths.
2.2 High confidence does not guarantee correctness
To move towards trustworthy AI in risk-sensitive domains such as medicine, it is crucial to develop systems that are not only correct but also confident and well-calibrated in their outputs (Gal et al., 2016; Lyu et al., 2024). Confidence, in practical terms, implies a higher degree of certainty in predictions. To assess this, we computed an entropy-based measure derived directly from model outputs rather than using a subjective self-assessment by LLMs (Jungo et al., 2020; Lyu et al., 2024; Wimmer et al., 2023; Zhou, 2020). Specifically, we compute normalized Shannon’s entropy () (Shannon, 1948) using the likelihood estimates for different SOZs and derive a confidence score () that ranges from 0 (lowest confidence) to 1 (highest confidence, see Eq. 5 in Methods). Across all models, confidence scores were lowest in the zero-shot condition and improved consistently with prompt engineering ( Figure 2(b)). For example, providing multiple solution examples in the few-shot condition increased the average model confidence by 13. 75% compared to zero-shot, while task-specific reasoning crafted by an epileptologist in the FS-CoT condition resulted in a substantial 21% increase in confidence, indicating that domain-specific demonstrations on task reasoning can significantly enhance the certainty of model predictions. Finally, as expected models exhibited the highest overall confidence for self-consistency (SC) with 35.25% improvement as this prompt style is intended to reduce stochastic variability by aggregating predictions across multiple reasoning paths. Assessing model confidence also requires evaluating calibration—that is how well model’s predicted probabilities align with its actual correctness. Even highly accurate models can produce probability estimates that do not reliably reflect their true likelihood of being correct (Guo et al., 2017). Calibration can be quantified using the Brier score loss (Brier, 1950; Niculescu-Mizil & Caruana, 2005), where lower scores indicate better alignment. As shown in Figure 2(c), models exhibit a high variance in calibration in the zero-shot condition. This markedly improves for all models with more refined prompting techniques. Similar to the effect we observed when assessing model confidence through entropy, FS-CoT and SC are most efficient for aligning predicted probabilities more closely with actual accuracy. Notably, GPT-4 shows the best calibration, even in the zero-shot condition.
Figure 2(d) provides a summary of model performance across three key dimensions (average across prompt styles): accuracy (F1 score), confidence, and calibration. Larger enclosed areas indicate models that better balance these factors, highlighting trade-offs between predictive accuracy and reliability. Given the medical domain’s high-risk nature, the ideal model is one that not only achieves strong performance but also maintains well-calibrated confidence, reducing the likelihood of overly certain but incorrect decisions. Overall, two models - GPT-4 and Mixtral-8x7B - strike the best balance between all factors in our task. We therefore consider these two models for a more detailed investigation, including reasoning analysis and experimental manipulations such as in-context impersonation, narrative length effects, and cross-linguistic comparisons.
2.3 Evaluating Clinical Reasoning and Source Attribution
Thus far, our results demonstrate that LLMs can effectively map unstructured seizure descriptions to seizure onset zones (SOZs) in the brain. To explore how LLMs arrive at their decisions, we assessed the reasoning abilities of the two best performing models, GPT-4 and Mixtral-8x7B on a randomly selected subset of 81 chain-of-thought (CoT) responses using a clinical evaluation (see Extended Data Figure 2 and Methods for details). Following Med-PaLM (Singhal et al., 2023a) and Liévin et al.’s (Liévin et al., 2024) protocol, we evaluated our models’ reasoning outputs using three categories: correct/complete, somewhat correct/complete, and incorrect/incomplete. Additionally, each output is assessed for the proportion of correct and incorrect statements along three dimensions: (i) comprehension, (ii) knowledge recall, and (iii) logical reasoning. A representative example output, where both models correctly identify the seizure onset zone (SOZ), along with the annotations, is shown in Figure 3(a). For the semiology “right upper limb ballistic-like movement”, both models demonstrate accurate comprehension, knowledge recall and logical reasoning (yellow, blue and green). However, in contrast to GPT-4, Mixtral-8x7B misinterprets the associated hemisphere leading to an error (see pink text: “right frontal lobe”). Additionally, Mixtral includes incorrect supporting scientific evidence, while GPT-4 cites a well-fitting and existing paper. These differences are well reflected in our summary analysis shown in Figure 3b-d. GPT-4 significantly outperformed Mixtral in both correctness (56.79% vs. 29.63%; z-test for proportions, p <0.05) and completeness (65.00% vs. 34.57%; z-test for proportions, p <0.05). Mixtral’s outputs were more often rated as “somewhat correct” and “somewhat complete” ((43.21%, 33.33% respectively). When examining specific dimensions ( Figure 3(c)), GPT-4 made fewer errors for knowledge recall than Mixtral (17.28% vs. 43.21%; z-test for proportions, p <0.05). However, the strongest differences were observed for logical reasoning: GPT-4 maintained a significantly higher score compared to Mixtral (98.77% vs. 80.25%; z-test for proportions, p <0.05). Similarly, Mixtral’s incorrect reasoning rate was almost twice that of GPT-4 (38.27% vs. 20.99%; z-test for proportions, p <0.05). Additionally, Figure 3(d) demonstrates GPT-4’s superior citation accuracy (76.54% for GPT-4 vs. 19.13% for Mixtral), indicating GPT-4’s enhanced ability to provide evidence-based sources for its decision.
2.4 Factors influencing LLM performance in seizure diagnostics
2.4.1 Symptom description length
Symptom descriptions vary in length—while longer descriptions may offer more information, they can also introduce irrelevant or contradictory details. To assess how this affects LLM-based diagnostic prediction, we categorized descriptions into six bins - based on word count and calculated the corresponding F1 scores, reporting the mean for each bin across all prompt strategies (refer Figure 4(a)). Interestingly, both models exhibit a distinctive U-shaped pattern, where predictions based on very short symptom descriptions initially achieve high performance, performance then decreases with increasing description length, and ultimately, observe the highest performance for the longest symptom descriptions. A Kruskal–Wallis test revealed statistically significant differences between the bins (p<0.001 for both GPT-4 and Mixtral-8x7B). We validated this pattern against surrogate data obtained by randomly sampling semiologies from the original data, preserving the bin sizes. We then recomputed the F1 scores for these shuffled bins. Kolmogorov-Smirnov test (p<0.001) between the original and shuffled data reveals that this U-shaped relationship represents a genuine effect of description length rather than a statistical artifact.
2.4.2 Clinical in-context impersonation
Clinical decision-making requires context-specific expertise. While prior work has shown that LLMs can improve performance by adopting expert personas in non-medical tasks (Salewski et al., 2024), we investigated whether this extends to clinical diagnostics. We prompted models to adopt increasingly specialized medical roles—medical assistant, neurologist, expert epileptologist—alongside a baseline AI assistant. Each prompt is prefixed with “you are a {persona}” and evaluated in a zero-shot setting to avoid few-shot example confounds. For both GPT-4 and Mixtral-8x7B, performance, and confidence improved with increasing domain specificity, with the “expert epileptologist” yielding the highest scores ( Figure 4(b)). GPT-4 showed a substantial performance gain of 13.68%, while Mixtral achieved a smaller but notable improvement (4.47%). Confidence scores also increased with more specialized personas (GPT-4: 9% and Mixtral: 8%). These results demonstrate that in-context persona adaptation enhances both performance and confidence in clinical tasks, with GPT-4 more effectively leveraging contextual cues.
2.4.3 Multilingual Performance
To assess the utility of LLMs for cross-lingual clinician-patient communication—particularly in medical tasks that depend entirely on verbal symptom descriptions—we compared performance across three language settings. In the baseline setting, both symptom descriptions and reasoning prompts were in English. In the mixed-language setting, symptom descriptions were in French, Spanish, or Chinese, while prompts remained in English. In the fully translated setting, both the clinical information and the reasoning instructions were presented entirely in French, Spanish, or Chinese, requiring full cross-lingual comprehension and reasoning. Our results (Figure Figure 4(c)) show that both models perform best when prompt and input are in English (En→En), likely reflecting their English-centric training. Interestingly, in the cross-language setting (non-English symptom, English prompt), performance only drops slightly (GPT-4: –1.91%, Mixtral: –2.41%, n.s.). However, in the same-language setting (non-English symptoms and prompt), Mixtral’s performance declines substantially by 8%, while GPT-4 remains stable (–1.4%, n.s.). This suggests both models can incorporate non-English input when anchored with English prompts, but especially Mixtral struggles in non-English contexts.
3 Discussion
Previous applications of LLMs in epilepsy have assessed general medical knowledge through structured Q&A formats using single models such as ChatGPT (Daungsupawong & Wiwanitkit, 2024; Holgate et al., 2024; Kim et al., 2024; Luo et al., 2024), while NLP approaches in epilepsy have used rule-based or supervised models trained for narrow tasks like seizure type classification and frequency extraction (Abeysinghe et al., 2025; Decker et al., 2022; Mora et al., 2024). Most studies evaluated factual recall, rather than diagnostic reasoning and lacked grounding in real-world patient data. In contrast, our work SemioLLM presents the first large-scale evaluation of 8 LLMs in diagnostic reasoning from over 1200 unstructured seizure descriptions. We observe that most LLMs can probabilistically infer seizure onset zones without structured input or domain-specific fine-tuning significantly above chance. Notably, GPT-4 and Mixtral-8x7B achieve performance comparable to a manual clinician-based assessment even under zero-shot conditions. Prompt engineering led to significant improvements in accuracy, confidence and calibration across all models - specifically clinician-guided chain-of-thought prompting led to the most substantial and consistent improvements. Our findings thus extend previous studies demonstrating the effectiveness of prompt engineering for structured medical Q&A (Singhal et al., 2023a), clinical name entity recognition (Hu et al., 2024), and medical text summarization (Sakai & Lam, 2025) to probabilistic reasoning from free-text symptom descriptions and especially shows the potential for integrating clinician expertise into foundation model–based clinical decision systems (Sonoda et al., 2025; Wu et al., 2023; Zou & Topol, 2025).
A key strength of LLMs is their capacity to generate explanatory reasoning alongside predictions—an important feature for transparency, interpretability, and trust in clinical decision support (Ghassemi et al., 2021). Adapting established evaluation protocols from Med-PaLM (Singhal et al., 2023a) and Liévin et al.(Liévin et al., 2024), we evaluated reasoning quality along with text comprehension, knowledge recall, and logical inference. Despite similar quantitative prediction performance, GPT-4’s reasoning outputs were more often rated as correct and complete in contrast to Mixtral, with a particularly strong advantage in logical inference. In contrast, Mixtral made reasoning errors in over a third of its outputs. Differences in citation accuracy were particularly strong (GPT-4: 76.54%, Mixtral: 19.13%), reflecting ongoing challenges in factual grounding and source attribution for generative models (Chen et al., 2023a; Ji et al., 2023). These results underscore that performance metrics alone may obscure reasoning deficiencies, findings that are known from studies testing general medical knowledge (Singhal et al., 2023b). Here, retrieval-augmented generation (RAG) (Lewis et al., 2020; Mialon et al., 2023) may help ground LLM reasoning in accurate, up-to-date knowledge, possibly improving reliability without retraining.
Our study also reveals critical factors influencing LLM performance on this task. First, similar to prior results in non-medical domains (Salewski et al., 2023), emulating an increasingly aligned clinical expert systematically improved both performance and confidence by 14% and 10% across models, respectively. Context-specific impersonation thus augments the ability of generalized language models to perform domain specific clinical tasks. Second, prediction performance varied with symptom description length, showing accuracy differences of up to 32%, where very short and highly detailed descriptions outperformed intermediate-length narratives. A possible explanation might be that brief descriptions closely match distinct canonical seizure features (e.g., “visual aura” for occipital onset (Williamson et al., 1992)), resembling the benefits of concise N-gram-based inputs in NLP tasks (Mahendra et al., 2021). Increasing description length, however, may introduce redundancies or contradicting evidence potentially degrading model performance (Liu et al., 2022). However, richly detailed but coherent input may offer enough structured context for LLMs to disambiguate and reason over complex clinical information. Ultimately these findings implicate that the level of detail as approximated through symptom description length might similarly influence diagnostic accuracy in LLMs and clinicians (Muayqil et al., 2018). Third, our evaluation showed that current LLMs can generalize across languages when anchored with English prompts, but also revealed limitations when all inputs are non-English. Both top-performing models, GPT-4 and Mixtral showed robust multilingual generalization when reasoning prompts were presented in English, even when symptom descriptions were provided in other languages such as French or Chinese, suggesting that they can integrate multilingual input when anchored with English-language instructions. In contrast, performance for one model (Mixtral) declined substantially when both prompts and input symptoms were non-English. This likely reflects the English-dominance of instruction tuning and pretraining corpora and points to limited cross-lingual generalization in current LLMs for the clinical domain, similar to limited multilingual performance in general-purpose NLP tasks (Hu et al., 2020). Targeted multilingual instruction tuning, particularly in clinical domains, may thus be necessary not only to ensure robust and inclusive model behavior but also equitable application in multilingual healthcare systems.
Unlike prior approaches that assess factual recall or classification, our framework combines quantitative metrics—correctness, confidence, and calibration — with qualitative expert annotation and evaluation of model reasoning. Importantly, LLMs can transform free-text clinical narratives into structured, actionable diagnostic inferences. Our work thus moves beyond basic knowledge verification and toward real-world clinical applicability. Importantly, our framework can be adapted to other medical domains. As a proof of concept, we provide example code and preliminary analyses in our GitHub repository in the domain of dermatology, linking skin anomaly descriptions to diagnostic classes.
Despite these strengths, our study has several limitations. Our analysis is restricted to a single dataset which includes only adult focal epilepsy cases. While the diagnostic task (localization of SOZ) may apply to pediatric patients, it is not suitable for generic, generalized seizures. Additionally, while the dataset includes cases from diverse sources, we did not have the necessary metadata to analyze the impact of demographic and cultural variations in patient descriptions. While our translation-based cross-lingual analysis suggests that the language of seizure descriptions can affect performance, a key limitation is that our original corpus was monolingual. It will therefore be important in future studies to quantify language-specific effects in culturally diverse seizure descriptions. We demonstrate in our public code repository that the framework can be adapted to other clinical domains such as dermatology, our systematic evaluation and prompt optimization were conducted specifically for epilepsy. Consequently, prompt engineering strategies may need to be re-evaluated and tailored for other specialties before clinical deployment. Finally, as a comprehensive manual annotation would require impractical amounts of clinician effort, our reasoning analysis focused on a representative subset, chosen to match the overall distribution of the data. While this approach provides practical insight, it may introduce a degree of sampling bias or subjectivity, and future studies with larger-scale or multi-annotator reasoning evaluations would further strengthen the conclusions.
4 Methods, Data and Code Availability
Details on the SemioLLM methodology, including the full process for prompt structuring, dataset preparation, and model evaluation protocols, are provided in the Methods section available online after the references. Additionally, we include extended data analyses and visualizations to support the findings presented in the main text.
The source code to reproduce the findings and figures are present at https://github.com/liebelab/semiollm
5 Acknowledgments
This work was supported by the Else Kröner Fresenius Foundation, Kolleg Clinbrain: Artificial Intelligence for Clinical Brain Research, the Clinician Scientist program of the Medical Faculty Tübingen funded by Deutsche Forschungsgemeinschaft (DFG, 493665037), the Priority Programme SPP 2241 - PN 520287829 (DFG), the Machine Learning Cluster of Excellence under Germany’s Excellence Strategy – EXC number 2064/1 PN 390727645, the Tübingen AI Center, the Collaborative Research Center 1233 “Robust Vision” funded by the DFG and ERC (853489-DEXIM). M.D. is a member of the International Max Planck Research School for Intelligent Systems Tübingen (IMPRS-IS). The authors thank Prof. Dr. Jakob H. Macke and his lab for discussion and feedback; Almut Sophia Koepke, Shyamgopal Karthik, Matthias Tangemann and Elisa Nguyen for comments on the manuscript.
6 Author Contribution
MD: Conceptualization, Methodology, Validation, Formal Analysis, Data Curation, Reasoning Study Design and Analysis, Visualization, Writing - Original Draft, Writing - Review and Editing. MJP: Methodology (design of clinical study form); Formal Analysis (compilation and evaluation of clinical responses and model source citations), Writing - Review and Editing. FP: Clinical evaluation, Writing - Review. ZA: Supervision, Funding Acquisition. SL: Conceptualization, Methodology, Clinical evaluation, Project Administration, Supervision, Funding Acquisition, Writing - Review and Editing.